OpenStudy (anonymous):

2-(9K-28)=6-15K PLEASE HELP!!!!!!!!!

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

actually i think you can distribute

OpenStudy (anonymous):

LOL I DONT KNOW WHICH ONE TO DO FIRST LOL

OpenStudy (anonymous):

2-(9K-28)=6-15K 2-9k+28=6-15k 30-9k=6-15k 30-6=15k+9k 24=24k k=1

OpenStudy (anonymous):

THANKX EYAD ARE U ARAB

OpenStudy (anonymous):

yw....

OpenStudy (anonymous):

ROFL DUDE ,i missed that ... and ya i'am kinda Arab ,Right guess Though :) and btw Yw = Your welcome = you are welcome ,what did u though it ?

OpenStudy (lalaly):

DOOOOOOOO YOUUUUUUUUU UNDERSTAND ARABICCCCCC!!!!!!!!!!!!!

OpenStudy (anonymous):

I FREAKIN ROFLEEED @lalaly

OpenStudy (lalaly):

hehe ma byeeeju ella bl3en el 7amra:P

OpenStudy (radar):

Does the answer k=1 really check out??

OpenStudy (anonymous):

LOL LALAY I DO I SOORY ILL CHAT LATER A HAVE BIG EXAM TOMOOROW I STUDY

OpenStudy (anonymous):

I DONT KNOW I RADER

OpenStudy (anonymous):

@radar, no it does not. K=-4

OpenStudy (radar):

2-(9(1)-28)=6-15(1) Needs further investigation..... 2-9+28=-9 21=-9 OhOh

OpenStudy (anonymous):

WAIT I THOUGHT IT WAS K=1 ROBTOBEY???!!! U PLEASE EXPLAIN

OpenStudy (radar):

We all can make mistakes @TAHA123.

OpenStudy (anonymous):

SO IT IS K=1

OpenStudy (anonymous):

ITS OKAY I UNDERSTAND

OpenStudy (radar):

As you can see 1 does not work, 21 does not equal -9

OpenStudy (radar):

No Eyad erred.

OpenStudy (anonymous):

WAIT I GOT IT DOES IT EQUAL -4

OpenStudy (radar):

If you will examine the fourth line of Eyad work, you will see the error.

OpenStudy (anonymous):

\[2-9k+28\text{ = }6-15k\]\[15k-9k\text{ = }6-30\]\[6 k=-24 \]\[K=-4 \]

OpenStudy (anonymous):

THANK VERY MUCH U HELP ME A LOT

OpenStudy (anonymous):

@TAHA123 Please don't write everything in capital letters ;)

OpenStudy (anonymous):

YYYYYYYYY